# Timer-XL: Long-Context Transformers for Unified Time Series Forecasting

### v3

@@ 4.1 Univariate Time Series Forecasting
Due to the insufficient dataset length when extending contexts in univariate datasets (Makridakis et al., 2020), we adopt wide-recognized benchmarks from Liu et al. (2023). Although these datasets are originally multivariate, they aim to be predicted in a univariate approach with the implementation of channel independence. Different from the previous long-term forecasting setting, we focus on reliable prediction based on a long context. Therefore, we fix the prediction horizon and increase the lookback length to monthly and yearly levels. We also establish a long-context univariate benchmark based on the challenging 40-year ECMWF Reanalysis v5 dataset (Hersbach et al., 2020), where yearly contexts are adopted to predict the land-surface temperature of a single site (ERA5-S).
@@ Results
As shown in Figure 3, the accuracy of univariate prediction can generally be improved by extending the daily context to monthly. We draw a similar conclusion on ERA5 (Table 14), where extending the context consistently helps in the specific model architecture. Empirically, Timer-XL, the generative Transformer, outperforms state-of-the-art encoder-only Transformer and linear forecaster in excessively long contexts. Further, we conduct representation analysis in Appendix E.4, revealing that Timer-XL is proficient at adaptively selecting information in vast observations, and thus achieves breakthrough performance. It is also noteworthy that the performance of monthly and yearly contexts improves slowly and deteriorates, which may stem from increased noise and training difficulty inherent in data, which leaves a future direction to improve the context efficiency. Table 1 provides results on ERA5-S. Timer-XL consistently outperforms PatchTST on all sites, which can be credited to the maintenance of causality and token-wise supervision in the decoder-only architecture.
Analysis
Furthermore, we analyze the widespread non-stationary challenge in univariate tasks. It is commonly tackled by normalization (Kim et al., 2021) that greatly improves Transformer performance in previous benchmarks. However, we find it may be caused by the insufficient time span and training samples in these datasets. Thus, normalization enriches training samples by aligning time series with different means and variances to the same distribution. Instead, it makes Transformer constrained on the temporal variation within windows, preventing them from learning variations among windows and resulting in oversmooth predictions and failures in long contexts. In Table 1 and Table 15, we evaluate the performance on ERA5 and widely-acknowledged datasets, which validates the claim that generative Transformers can achieve better results even without instance normalization.
Table 1: Univariate forecasting (input-$3072$-pred-$96$) of ERA5-S ($40$ years), encompassing $117$k time points in each station. We evaluate PatchTST and Timer-XL with and without normalization (Kim et al., 2021). + Norm. indicates using the normalization. We train one model for each site separately.
@@ 4.2 Multivariate Time Series Forecasting
We follow iTransformer (Liu et al., 2023) to evaluate multivariate forecasting performance. Toward a one-for-all forecaster, we also evaluate rolling forecast performance, that is, we trained one model for all prediction horizons by integrating the previous prediction into the lookback window in the next iteration. We further establish long-context multivariate forecasting benchmarks: ERA5 multi-station land-surface temperature prediction (ERA5-MS), and the global temperature and wind speed forecasting challenge (GTWSF) (Wu et al., 2023), to learn complex temporal dynamics and variable correlations with sufficient training samples.
@@ Results
Analysis
Patching (Nie et al., 2022) has been demonstrated as an effective tokenization approach for the time series modality, leading to the boom of Transformers in supervised deep forecasters and large time series models. To better cope with multivariate time series forecasting, we compared these Transformers on ERA5-MS to answer the following questions: (1) whether to conduct explicit inter-series modeling or not (channel independence) and (2) whether to use decoder-only or encoder-only Transformers. The combination presents four typical Transformers in Table 4, which shows that Timer-XL combines the advantages of explicit inter-series modeling and the decoder-only architecture, which is suitable for multivariate time series forecasting with arbitrary prediction horizons.
@@
4.4 Pre-Trained Time Series Transformers
@@
Pre-training enriches time-series Transformers with generalizable forecasting capabilities. The outcome large time series model can better cope with widespread challenges of few-shot and zero-shot forecasting. To scale Timer-XL as large models, we conduct pre-training on UTSD (Liu et al., 2024c) and LOTSA (Woo et al., 2024) for zero-shot forecasters. We also curate a large-scale multivariate dataset (ERA5-Large: 40 years and 4920 stations) for a domain-specific model. In this task, we evaluate in-dataset generalization performance of PatchTST and Timer-XL: pre-training on $80$% stations and $80$% time span and then forecast on the remaining stations (variable generalization), time span (temporal generalization), and the cross-time and cross-station split. To evaluate the benefit of pre-training with longer context, we compare the zero-shot performance of Timer (2024c) and Timer-XL by pre-training on UTSD (1B time points), where the context length of Timer-XL is increased from $672$ to $2880$. To establish a fair zero-shot benchmark, we pre-train Timer-XL on LOTSA (27B observations) such that baseline models are built upon a comparable pre-training scale.
@@ Results
We provide in-dataset generalization results of ERA5-Large in the middle of Figure 5 (a). Timer-XL achieves better results than PatchTST in all cases, supporting that decoder-only architecture has stronger generalization performance. Figure 5 (b) compares zero-shot performance of two generative Transformers pre-trained on the same UTSD, where Timer-XL outperforms previous Timer on all benchmark datasets, validating that long-context pre-training empowers large time series models. In Table 6, we provide a comprehensive zero-shot evaluation under a comparable pre-training scale and model size, where Timer-XL achieves notable performance with better sample efficiency. The versatility and scalability make it a promising backbone of foundation models.
@@
  • $\ast$ Traffic is not evaluated here because it is generally included in the pre-training datasets of large models.
@@ 4.5 Model Analysis
To evaluate the model efficiency of Timer-XL with respect to the context length, it is essential to recognize the distinct characteristics of time series data compared to 1D sequences. Unlike natural language, the time series modality is characterized by the variable number $N$ and the input length. We adopt two representative multivariate datasets with different $N$, and provide the memory footprint and training speed under gradually prolonged input. We evaluate typical approaches to handle multivariate series: (1) Timer-XL and Moiria that adopt channel dependence; (2) Timer that adopts channel independence. Intuitively, the complexity of the first type is $\mathcal{O}(N^{2}T^{2})$ while the complexity of self-attention under channel independence is $\mathcal{O}(NT^{2})$. However, results shown in Figure 6 reveal that the measured cost of Timer-XL is much less than $N$ times of Timer.
Since the previous analysis of model efficiency on time-series Transformer predominantly focuses on the self-attention on 1D time series, we initially present a theoretical derivation of the computational complexity of Transformers on 2D time series, including the parameter counts, memory footprint, and FLOPs in Table 7. We find that other parts of Transformers, such as feed-forward network, have a complexity of $\mathcal{O}(NT)$ no matter which approach is adopted to handle multivariate time series. They also cause the dominant overhead in existing benchmarks, since their context length is not large enough, confirming our empirical results. Further, we introduce FlashAttention (Dao et al., 2022) to reduce the memory footprint and training speed, which is computationally equivalent and reduces the overall memory footprint of Timer-XL to $\mathcal{O}(NT)$ without affecting performance.
@@
Figure 7: Visualization of TimeAttention. It is from the first sample of a length $672$ in the test split of Traffic. We visualize the last $10$ variables and each contains $7$ tokens. We present auto-correlation function plot. Auto-correlation can be reflected by the distribution of the attention (bottom right). We average TimeAttention across sub-blocks, which can reveal Pearson correlations (upper right).
@@ 5 Conclusion and Future Work
Grounded in the principles of forecasting, we highlight the urgency to extend the context length in the time series field. To facilitate long-context forecasters on diverse tasks, we propose multivariate next token prediction, a novel paradigm to predict 1D and 2D time series with covariates. We present Timer-XL enhanced by TimeAttention as an extra-long version of generative time-series Transformers. It simultaneously captures temporal dynamics and variable correlations by causal self-attention. In addition to achieving state-of-the-art performance on extensive datasets, we establish challenging benchmarks for long-context forecasting. Further, by pre-training on large-scale heterogeneous time series, Timer-XL demonstrates significant generalization capabilities as a one-for-all large model. In the future, we will improve the context utilization and computational efficiency.
@@ A.1 Setups
Metric Type Count / Complexity
FLOPs Channel Independence (Nie et al., 2022) $\mathcal{O}(PDNT+L(D+H)NT^{2}+(2+\alpha)LD^{2}NT)$
@@
Channel Dependence (Ours) $\mathcal{O}(PDNT+L(D+H)N^{2}T^{2}+(2+\alpha)LD^{2}NT)$
Parameters Flatten head (Nie et al., 2022) $(4+2\alpha)LD^{2}+4LD+(1+T)PD$
Token-wise projector (Liu et al., 2024c) $(4+2\alpha)LD^{2}+4LD+2PD$
Memory Self-Attention (Vaswani et al., 2017) $4(D+P)NT+(32+8\alpha)LDNT+4LHN^{2}T^{2}$
Footprint FlashAttention (Dao et al., 2022) $4(D+P)NT+(32+8\alpha)LDNT$
  • $\ast$ $L$ is the block number of Transformers. $D$ is the dimension of embeddings (the hidden dimension of FFN $D_{\text{ff}}$ is set as $\alpha D$). $H$ is the head number and the dimension of query, key, and value $d_{k}=D/H$. The overhead is to train on a multivariate time series ($N$-variables and $TP$ time points) and predict the next patch with the length $P$.
@@ B.1 Datasets
  • ERA5-MS: Each univariate series of ERA5-S provides partial observations governed by the spatio–temporal global weather system. Since discovering the global spatio-temporal correlations presents a fundamental challenge in meteorology, we convert ERA5-S into ERA5-MS by using seven subsets as a challenging multivariate forecasting benchmark. Based on the average results in Tables 1 and 4, we can validate the existence of multi-station correlations among selected stations, which have enhanced the average prediction accuracy.
@@
Tasks Dataset Dim. Prediction Setting Dataset Length
@@ Appendix C Hyperparameter Sensitivity
We evaluate the hyperparameter sensitivity of Timer-XL on the ERA5-MS benchmark, as illustrated in Figure 8, concerning the following factors: the number of layers $L$, the patch size $P$, and the lookback length during inference. Our findings indicate that performance of Timer-XL generally improves with increases with $L$, suggesting that Timer-XL is a scalable deep forecaster. Furthermore, our analysis of the influence of $P$ reveals that the optimal patch size is generally close to the predicted length, since it avoid multi-step error accumulations. Toward better long-term forecasting performance, it leaves a future improvement to adopt different patch sizes of input and output tokens. Finally, we investigate the impact of input length during inference. We discover that the optimal lookback length of during is not necessarily the length during training. Given that generative Transformers can accommodate inference inputs shorter than those used during training, this finding is noteworthy and indicates the potential to improve the performance of generative Transformers.
@@ E.2 Full Result of Zero-Shot Forecasting
Table 12: Full results of zero-shot forecasting (pred-$\{96,192,336,720\}$). We add new baseline large models, such as Time-MoE (Shi et al., 2024). Our model size is comparable with $\textbf{Moirai}_{\textit{Base}}$, which is pre-trained on UTSD (Liu et al., 2024c) and LOTSA (Woo et al., 2024).
@@
  • $\ast$ Traffic from Wu et al. (2022) is generally used during the pre-training of large models and thus not evaluated here.
@@ E.4 Supplementary Results of Long-Context Forecasting
Section 4.1 has discussed instance normalization (Kim et al., 2021). It generally improves the performance of the previous encoder-only Transformers but leads to special problems in generative Transformers (e.g., unmatched statistics in multi-step autoregression). However, it is indicative that Timer-XL without ReVIN can achieve competitive performance on well-acknowledged benchmarks in Table 15, while the performance of PatchTST may heavily rely on this normalization.
@@ E.5 Illustration of TimeAttention
Although the formulation to generalize from 1D sequences to multivariate time series is straightforward, Timer-XL is built on a generative Transformer, an underexploited backbone among current time series models. As shown in Figure 12, challenges lie in capturing fine-grained dependencies between all variables in the patch level, while maintaining temporal causality in multiple sequences. Technically, we introduce the masking formulation, whose key lies in the grouped causality of flattened 2D sequences. We derive it based on the Kronecker Product, which disentangles the large attention map into formalizable temporal and variable dependencies. It can be naturally extended to covariates or pre-defined variable dependencies, which may inspire a lot of future explorations.
@@ Appendix F Limitations
As a generative Transformer, Timer-XL necessitates iterative generation for long-term forecasting, which may lead to error accumulation and inflexibility in the output length. In the future, we plan to incorporate multi-resolution patches for input and output series. Furthermore, given that Timer-XL explicitly captures fine-grained token dependencies, there remains significant potential to reduce the complexity of TimeAttention, particularly in high-dimensional and lengthy time series. Finally, we will investigate the factors contributing to the stagnation of Transformer performance in extremely long contexts, and seek insights in the time series modality to improve context efficiency.

### v4

@@ 4.1 Univariate Time Series Forecasting
Due to the insufficient dataset length when extending contexts in univariate datasets (Makridakis et al., 2020), we adopt multivariate datasets from Liu et al. (2023). Although these datasets are originally multivariate, they aim to be predicted in a univariate approach with the implementation of channel independence. Different from the previous long-term forecasting setting, we focus on reliable prediction based on a long context. Therefore, we fix the prediction horizon and increase the lookback length to monthly and yearly levels. We also establish a long-context univariate benchmark based on the challenging 40-year ECMWF Reanalysis v5 dataset (Hersbach et al., 2020), where yearly contexts are adopted to predict the land-surface temperature of a single site (ERA5-S).
@@ Results
As shown in Figure 3, the accuracy of univariate prediction can generally be improved by extending the daily context to monthly. We draw a similar conclusion on ERA5 (Table 15), where extending the context consistently helps in the specific model architecture. Notably, Timer-XL with decoder-only architecture outperforms encoder-only Transformer and linear forecaster in excessively long contexts. Further, we conduct representation analysis in Appendix E.4, revealing that Timer-XL is proficient at adaptively selecting information in vast observations and thus achieves breakthrough performance. It is also noteworthy that the performance of monthly and yearly contexts improves slowly and deteriorates, which may stem from increased noise and training difficulty inherent in data, which leaves a future direction to improve the context efficiency. Table 2 provides results on ERA5-S. Timer-XL consistently outperforms PatchTST on all sites, which can be credited to the maintenance of causality and token-wise supervision in the decoder-only architecture.
Non-stationary Forecasting
We delve into widespread non-stationarity in univariate tasks. It is commonly tackled by normalization (Kim et al., 2021) that greatly improves Transformer performance in previous benchmarks. However, we find it may be caused by the insufficient time span and training samples in these datasets. While normalization simplifies learning by aligning series with different means and variances to the same distribution, it limits the model capacity of Transformers, preventing them from learning variations among windows. The by-product can be mode collapse and oversmooth predictions. In Table 2 and Table 16, we evaluate the performance on ERA5 and datasets from Wu et al. (2022), which validates that Timer-XL can achieve better results even without instance normalization.
Table 2: Univariate forecasting (input-$3072$-pred-$96$) of ERA5-S, encompassing $117$k time points in each station ($40$-years). We evaluate PatchTST and Timer-XL with and without normalization (Kim et al., 2021). + Norm. indicates using the normalization. We train one model for each site separately.
@@ 4.2 Multivariate Time Series Forecasting
We follow iTransformer (Liu et al., 2023) to evaluate multivariate forecasting performance. Toward a one-for-all forecaster, we evaluate performance of rolling forecast, that is, we trained one model for all prediction horizons by integrating the previous prediction into the lookback window in the next iteration. We further establish long-context multivariate forecasting benchmarks: ERA5 multi-station land-surface temperature prediction (ERA5-MS), and the global temperature and wind speed forecasting challenge (GTWSF) (Wu et al., 2023), to learn complex temporal dynamics and variable correlations with sufficient training samples.
@@ Results
Ablation Study
Patching (Nie et al., 2022) has been demonstrated as an effective tokenization approach for time series, leading to the boom of Transformers in supervised deep forecasters and large time series models. To better cope with multivariate time series forecasting, we compared typical models on real-world benchmarks to address key questions: (1) whether to conduct explicit inter-series modeling or not (channel independence) and (2) whether to use decoder-only or encoder-only Transformers. The combination presents four Transformers in Table 5, which shows that Timer-XL combines the advantages of explicit inter-series modeling and the decoder-only architecture, which is suitable for multivariate time series forecasting with sufficient training samples.
@@
4.4 Pre-Trained Time-Series Transformers
@@
Pre-training enriches time-series Transformers with generalizable forecasting capabilities. The outcome large time series model can cope with widespread challenges of few-shot and zero-shot forecasting. In this section, we conduct univariate pre-training on UTSD (Liu et al., 2024c) and LOTSA (Woo et al., 2024) and evaluate zero-shot performance on benchmarks from Wu et al. (2022). We further conduct large-scale multivariate pre-training on our ERA5-Large dataset, which spans 40 years and encompasses 4920 stations. Subsequently, we evaluate three types of generalization results comparing PatchTST (encoder-only Transformer) and Timer-XL (decoder-only Transformer): pre-training on $80$% stations and $80$% time span and then forecast on the remaining stations (variable generalization), remaining time span (temporal generalization), and remaining split of time span and stations (variable and temporal generalization). To evaluate the benefit of pre-training with longer context, we compare the zero-shot performance of Timer (2024c) and Timer-XL, where the context length of pre-training is increased from $1440$ to $2880$.
@@ Results
We compare generalization performance on ERA5-Large in the middle of Figure 5 (a). Timer-XL achieves better results than PatchTST in all cases, revealing that decoder-only architecture has stronger generalization capability. Figure 5 (b) compares zero-shot performance of two pre-trained Transformers with different context lengths, where Timer-XL outperforms previous Timer on all benchmark datasets, validating that long-context pre-training enhances large time series models. In Table 7, we provide a comprehensive zero-shot evaluation under a comparable pre-training scale and model size, where Timer-XL achieves notable performance with better sample efficiency. The versatility and scalability make it a promising backbone of foundation models.
@@
5
1
10
0
1
2
0
0
0
0
2
  • $\ast$ Traffic from (PEMS, ) is generally used during the pre-training of large models and thus not evaluated here.
@@ 4.5 Model Analysis
To evaluate the model efficiency of Timer-XL with respect to the context length, it is essential to recognize the distinct characteristics of time series data compared to 1D sequences. Unlike natural language, the time series modality is characterized by the variable number $N$ and the input length. We adopt two representative multivariate datasets with different $N$, and provide the memory footprint and training speed under gradually prolonged input. We evaluate typical approaches to handle multivariate series: (1) Timer-XL and Moiria that adopt channel dependence; (2) Timer that adopts channel independence. Intuitively, the complexity of the first type is $\mathcal{O}(N^{2}T^{2})$ while the complexity of self-attention under channel independence is $\mathcal{O}(NT^{2})$. However, results shown in Figure 6 reveal that measured overheads of Timer-XL is much less than $N$ times of Timer.
Since the previous analysis of model efficiency on time-series Transformer predominantly focuses on self-attention on 1D time series, we initially present a theoretical derivation of the computational complexity of Transformers on 2D time series, including the parameter counts, memory footprint, and FLOPs in Table 8. We find that other parts of Transformers, such as feed-forward network, have a complexity of $\mathcal{O}(NT)$ no matter which approach is adopted to handle multivariate time series. They also account for dominant overheads in existing benchmarks since the context length is not large enough, confirming our empirical results. Further, we introduce FlashAttention (Dao et al., 2022) to improve the model efficiency, which is computationally equivalent and reduces the overall memory footprint of Timer-XL to $\mathcal{O}(NT)$ without affecting performance.
@@
Figure 7: Visualization of TimeAttention. It is from the first sample of a length $672$ in the test split of Traffic. We visualize the last $10$ variables with each contains $7$ tokens. We present auto-correlation function plot. Auto-correlation can be reflected by the distribution of attention scores (bottom right). We average TimeAttention across sub-blocks, which indicates Pearson correlations (upper right).
@@ 5 Conclusion and Future Work
In this paper, we emphasize the efficacy of causal Transformers in the forecasting of long-context time series. To facilitate long-context Transformers on diverse tasks, we propose multivariate next token prediction, a novel paradigm to predict multidimensional series with covariates. We present Timer-XL enhanced by TimeAttention as an extra-long version of pre-trained time-series Transformers. It simultaneously captures temporal dynamics and variable correlations by enhanced self-attention. In addition to achieving state-of-the-art performance on extensive benchmarks, we establish challenging benchmarks for long-context forecasting. By pre-training on large-scale heterogeneous time series, Timer-XL demonstrates notable zero-shot performance as a large time-series model. In the future, we will improve computational efficiency and build large domain-specific models with Timer-XL.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (U2342217 and 62021002), the BNRist Project, and the National Engineering Research Center for Big Data Software.
@@ A.1 Setups
Metric Type Count Complexity
FLOPs Channel Independence $12\big{(}PDNT+L(D+H)NT^{2}+(2+\alpha)LD^{2}NT\big{)}$ $\mathcal{O}\big{(}LDNT(D+T)\big{)}$
@@
Channel Dependence $12\big{(}PDNT+L(D+H)N^{2}T^{2}+(2+\alpha)LD^{2}NT\big{)}$ $\mathcal{O}\big{(}LDNT(D+NT)\big{)}$
Parameters Encoder-Only $(4+2\alpha)LD^{2}+4LD+(1+T)PD$ $\mathcal{O}(LD^{2})$
Decoder-Only $(4+2\alpha)LD^{2}+4LD+2PD$ $\mathcal{O}(LD^{2})$
Memory Self-Attention $4(D+P)NT+(32+8\alpha)LDNT+4LHN^{2}T^{2}$ $\mathcal{O}\big{(}LHN^{2}T^{2}\big{)}$
Footprint FlashAttention $4(D+P)NT+(32+8\alpha)LDNT$ $\mathcal{O}\big{(}LDNT\big{)}$
  • $\ast$ $L$ is the block number of Transformers. $D$ is the dimension of embeddings (the hidden dimension of FFN $D_{\text{ff}}$ is set as $\alpha D$). $H$ is the head number and the dimension of query, key, and value $d_{k}=D/H$. The overhead is to train on a multivariate time series ($N$-variables and $TP$ time points) with patch token length $P$ and context length $T$. Set $N=1$ for training on univariate time series.
@@ B.1 Datasets
  • ERA5-MS: Each univariate series of ERA5-S provides partial observations governed by the spatio–temporal global weather system. Since discovering the global spatio-temporal correlations presents a fundamental challenge in meteorology, we convert ERA5-S into ERA5-MS by using seven subsets as a challenging multivariate forecasting benchmark. Based on the average results in Tables 2 and 5, we can validate the existence of multi-station correlations among selected stations, which have enhanced the average prediction accuracy.
@@
Tasks Dataset Dim. Training Setting Dataset Length
@@ Appendix C Hyperparameter Sensitivity
We evaluate the hyperparameter sensitivity of Timer-XL on the ERA5-MS benchmark, as illustrated in Figure 8, concerning the following factors: the number of layers $L$, the patch size $P$, and the lookback length during inference. Our findings indicate that performance of Timer-XL generally improves with increases with $L$, suggesting that Timer-XL is a scalable deep forecaster. Furthermore, our analysis of the influence of $P$ reveals that the optimal patch size is generally close to the predicted length, since it avoid multi-step error accumulations. Toward better long-term forecasting performance, it leaves a future improvement to adopt different patch sizes of input and output tokens. Finally, we investigate the impact of input length during inference. We discover that the optimal lookback length of during is not necessarily the length during training. Given that decoder-only Transformers can accommodate inference inputs shorter than those used during training, this finding is noteworthy and indicates the potential to improve the performance.
@@ E.2 Full Result of Zero-Shot Forecasting
Table 13: Full results of zero-shot forecasting. A lower MSE or MAE indicates a better prediction. $1^{\text{st}}$ Count represents the number of wins achieved by a model under all prediction lengths and datasets.
@@
  • $\ast$ Traffic from (PEMS, ) is generally used during the pre-training of large models and thus not evaluated here.
@@ E.4 Supplementary Results of Long-Context Forecasting
Section 4.1 has discussed instance normalization (Kim et al., 2021). It generally improves the performance of the previous encoder-only Transformers but leads to special problems in decoder-only Transformers (e.g., unmatched statistics in multi-step autoregression). However, it is indicative that Timer-XL without ReVIN can achieve competitive performance on well-acknowledged benchmarks in Table 16, while the performance of PatchTST may heavily rely on this normalization.
@@ E.5 Illustration of TimeAttention
Although the formulation to generalize from 1D sequences to multivariate time series is straightforward, Timer-XL is built on a decoder-only Transformer, an underexploited backbone among current time series models. As shown in Figure 12, challenges lie in capturing fine-grained dependencies between all variables in the patch level, while maintaining temporal causality in multiple sequences. Technically, we introduce the masking formulation, whose key lies in the grouped causality of flattened 2D sequences. We derive it based on the Kronecker Product, which disentangles the large attention map into formalizable temporal and variable dependencies. It can be naturally extended to covariates or pre-defined variable dependencies, which may inspire a lot of future explorations.
@@ Appendix F Limitations
Timer-XL is a unified model for time series forecasting. It can be used for task-specific training or scalable pre-training, handling varying-length and multivariate time series. As an autoregressive model, Timer-XL necessitates iterative generation for long-term forecasting, which may lead to error accumulation and inflexibility in the output length. In the future, we plan to incorporate multi-resolution patches for input and output series. Furthermore, given that Timer-XL explicitly captures fine-grained token dependencies, there remains significant potential to reduce the complexity of TimeAttention, particularly in high-dimensional and lengthy time series. Finally, we will investigate the factors contributing to the stagnation of Transformer performance in extremely long contexts, and seek insights in the time series modality to improve context efficiency.